OpenStudy (anonymous):

Alicia Finished the race 3.75 seconds before Kaye. Donna finished the race 5.09 seconds after Carmen. Carmen finished the race 10 seconds before Kaye. Anders finished 4.89 seconds after Kaye. how many seconds earlier did Carmen finish than Anders??

OpenStudy (anonymous):

A=K-3.75 D=C+5.09 C=K-10 An=k-4.89

OpenStudy (anonymous):

thanks your awesome

OpenStudy (anonymous):

That all you wanted? Are you sure?

OpenStudy (anonymous):

Those are just the formulas

OpenStudy (anonymous):

well i don't need all the other extraneous info about kaye and donna right?

OpenStudy (anonymous):

I'm new to this website

OpenStudy (anonymous):

If you make a chart of who finishes first, you will get the answer

OpenStudy (anonymous):



OpenStudy (anonymous):